OpenStudy (anonymous):

1/2(log base 5 x + log base 5 y)- 5log base 5 (x+8) Write as a single log and evaluate if possible.

OpenStudy (anonymous):

\[\frac12(\log_{5} x+\log_{5} y)-5\log_{5} (x+8)\]

OpenStudy (anonymous):

\[\frac12(\log_5(xy))-5\log_5(x+8)\]

OpenStudy (anonymous):

keep going....I need a single log expression

OpenStudy (anonymous):

\[\log_5\sqrt {xy}-\log_5(x+8)^5\]

OpenStudy (anonymous):

\[\log_5\left( \frac{\sqrt {xy}}{(x+8)^5}\right) \]

OpenStudy (anonymous):

I guess this is enough

OpenStudy (anonymous):

yeah...that was correct....thanks!!!

OpenStudy (anonymous):

np